# WaveCert: Formal Compiler Validation for Asynchronous Dataflow Programs

### v1

@@ 5. Confluence Checking using Linear Permission Tokens
After checking cut-bisimilarity between the LLVM program and the dataflow program on the canonical schedule, we still need to prove that this canonical schedule is general. Indeed, since the dataflow program is asynchronous and nondeterministic, other choices of schedule may lead to data races and inconsistent final states.
Therefore, in addition to finding a bisimulation on the canonical schedule, we want to prove a confluence property on the dataflow program:
Definition 0 ().
@@
In this section, we present a technique using what we call linear permission tokens to ensure that the dataflow program is confluent. Linear permission tokens are, in essence, a dynamic variant of the fractional permissions studied in the concurrent separation logic (fractional-perm, ; concurrent-separation-logic, ). In our case, their linear usage is dynamically enforced and the exact placement of permission tokens is resolved after symbolic execution.
@@
We associate each value present in a dataflow execution trace with a permission token in the form of $\texttt{write}\ l$, $\texttt{read}\ l$, or any formal sum of them, for any memory location $l$. These tokens are used by dataflow operators in a linear fashion: they cannot be replicated, and they are disjoint in the initial state. A write token can be split into $n$ many read tokens (for a predetermined parameter $n$), and inversely, $n$ of such read tokens can be combined to get back a write token. As shown later in this section, if there is a consistent assignment of these linear permission tokens, we can conclude that there are no data races possible in any schedule, and thus all final states have to be the same.
@@
In Section 5.1, we prove that the technique works in a bounded setting for a concrete trace $\tau$: if $\tau$ has a consistent assignment of permission tokens, then for any other trace $\tau^{\prime}$ bounded by $\tau$ (meaning that for each operator, $\tau^{\prime}$ executes it for an equal or smaller number of times than $\tau$), $\tau^{\prime}$ will converge to the final configuration of $\tau$.
@@
Then in Section 5.2, we explain how this bounded confluence check can be extended to a full confluence check for the entire dataflow program by using the cut points that we have placed for the dataflow program.
@@ 5.1. Confluence on Bounded Traces
In this section, we show that if we have a $n$-step trace $c_{1}\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}c_{n+1}$ (where $o_{1},\dots,o_{n}$ are the operators fired in each step) with a consistent assignment of permission tokens to each value, then for any other trace $c_{1}\to^{o^{\prime}_{1}}_{\mathrm{df}}\dots\to^{o^{\prime}_{m}}_{\mathrm{df}}c_{m+1}$ (not necessarily with a consistent assignment of permission tokens) such that the multiset $\{o^{\prime}_{1},\dots,o^{\prime}_{m}\}\subseteq\{o_{1},\dots,o_{n}\}$, we have $c_{m+1}\to^{*}_{\mathrm{df}}c_{n+1}$.
@@
  • $(P(L,k),0,+)$ is the term algebra of the formal sums $\texttt{read}\ l_{1}+\dots+\texttt{read}\ l_{n}$, with $l_{1},\dots,l_{n}\in L$, modulo the commutativity and associativity of $+$, and we require that for a location $l$, at most $k$ copies of $\texttt{read}\ l$ is present in the formal sum. We call $p+q$ the disjoint union of $p,q\in P(L,k)$. We also denote
@@
In Boyland’s terminology (fractional-perm, ), $P(L,k)$ is a finite token model of size $k$ of fractional permissions. For example, in $P(\{\texttt{A},\texttt{B}\},2)$, we have that
@@
Definition 0 ().
We say that the permissions $p_{1},\dots,p_{n}\in P(L,k)$ are disjoint iff for any $i\neq j$, $p_{i}+p_{j}$ is defined.
Now let us define an assignment of permission tokens to values in a configuration. Let $E$ denote the set of channel names. For a configuration $c\in\mathcal{C}_{\mathrm{df}}$ and a channel $e\in E$, let $\mathrm{Channel}(c,e)$ denote the state of the channel $e$ in $c$, which is a string over $V$ representing the values in the channel $e$.
@@
Let $c\in\mathcal{C}_{\mathrm{df}}$ be a dataflow configuration. A permission augmentation of $c$ is a partial map $t:E\times\mathbb{N}\to P(L,k)$ such that for any $e\in E$, $t(e,n)$ is defined iff $n\in\{1,\dots,|\mathrm{Channel}(e)|\}$. We call $(c,t)$ a permission-augmented configuration.
Intuitively, $t(e,n)$ is the permission we attach to the $n$-th value in the channel $e$. In the following, we define when such a permission augmentation is consistent (i.e., the use of permissions is linear).
@@
Let $c\to^{o}_{\mathrm{df}}c^{\prime}$ be a transition. Let $t,t^{\prime}$ be two permission augmentations for $c,c^{\prime}$, respectively. Let $p_{1},\dots,p_{n}$ be the permissions in $t$ attached to input values of $o$ and $q_{1},\dots,q_{m}$ be the permissions in $t^{\prime}$ attached to output values of $o$. We say $(c,t)\to^{o}_{\mathrm{df}}(c^{\prime},t^{\prime})$ is a consistent (permission-augmented) transition iff:
  • The $t$ and $t^{\prime}$ have the same assignment of permissions except for the channels changed by the operator $o$.
@@
Let $c_{1}\to_{\mathrm{df}}\dots\to_{\mathrm{df}}c_{n}$ be a trace. Let $t_{1},\dots,t_{n}$ be permission augmentations to $c_{1},\dots,c_{n}$, respectively. We say that $(c_{1},t_{1})\to_{\mathrm{df}}\dots\to_{\mathrm{df}}(c_{n},t_{n})$ is a consistent (permission-augmented) trace iff:
  • $(c_{1},t_{1})$ is consistent
  • $(c_{i},t_{i})\to_{\mathrm{df}}(c_{i+1},t_{i+1})$ is consistent for $i\in\{1,\dots,n\}$.
The following lemma allows us to swap the order of execution of operators in two consistent steps.
Lemma 8 ().
@@
Proof Sketech.
If one of $o_{1}$ and $o_{2}$ is neither load nor store, or if both of them are loads, the result will always be consistent.
If $o_{1}$ and $o_{2}$ are both stores, and they are the same node, then it must be the case that $c^{\prime}_{2}=c_{2}$ and we can use $t^{\prime}_{2}=t_{2}$.
Otherwise, assume both $o_{1}$ and $o_{2}$ are load/store operators, and they are different nodes in the dataflow graph. Since they can both fire at $c_{1}$, and they are different nodes, their input values must be attached with disjoint permissions. Therefore, $o_{1}$ and $o_{2}$ have no dependencies, and they will give the same result no matter in which order we execute them. ∎
@@
Practically, with this bounded confluence theorem, we can do bounded model checking for the confluence property: we perform symbolic execution to obtain a trace of symbolic configurations $c_{1}(\mathbf{x})\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}c_{n+1}(\mathbf{x})$, and then we can solve for a consistent augmentation $(t_{1},\dots,t_{n+1})$ for this trace using an SMT (z3, ) solver. If there is a solution, then it follows from Theorem 10 that any other trace (firing operators only in the multiset $\{o_{1},\dots,o_{n}\}$, counting multiplicity) would still converge to $c_{n+1}(\mathbf{x})$ in the end.
@@ 5.2. Full Confluence Checking using Cut Points
To extend the bounded confluence results in Section 5.1 to checking confluence for the entire dataflow program with unbounded traces, we need to finitely describe and check a consistent pattern of permission augmentations for all traces. For this purpose, we use the cut points selected in Section 4.
@@
Let $P^{1}_{\mathrm{df}}(\mathbf{x^{1}}),\dots,P^{n}_{\mathrm{df}}(\mathbf{x^{n}})$ be a list of cut points. During the bisimulation check in Section 4.2, we have verified that starting from any cut point $P^{i}_{\mathrm{df}}(\mathbf{x^{i}})$, any branch of symbolic execution reaches an instance of another cut point $P^{j}_{\mathrm{df}}(\mathbf{x^{j}})$. A consequence of this check is that all the symbolic branches from all cut points “cover” any complete trace (traces from an initial configuration to a final configuration). In other words, any complete concrete trace from the initial configuration is the concatenation of instances of symbolic branches.
@@
Thus, if we are able to find a consistent permission augmentation for the trace of each symbolic branch during the bisimulation check, then any concrete trace in $\to_{\mathrm{df\mid canon}}$ would have a consistent permission augmentation, which implies that any other schedule will converge to $\to_{\mathrm{df\mid canon}}$, and the confluence of the entire dataflow program follows.
Theorem 11 ().
If all symbolic dataflow branches produced in Section 4.2 have consistent permission augmentations, then $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ is confluent.

### v2

@@ 5. Confluence Checking using Linear Permission Tokens
After checking the equivalence between the LLVM and dataflow programs on a canonical dataflow schedule, we still need to prove that this canonical schedule is general. Indeed, since the dataflow program is asynchronous and nondeterministic, other choices of schedule may lead to data races and inconsistent final states.
Therefore, in addition to checking the equivalence on a canonical schedule, we want to prove a confluence property on the dataflow program:
Definition 0.
@@
If the pair of dataflow and LLVM programs satisfies both Definition 4 and Definition 1, then we can conclude the desired equivalence property of Definition 2.
Proposition 0.
If $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{llvm}},\mathcal{C}_{\mathrm{llvm}},\to_{\mathrm{llvm}})$ are equivalent on a canonical schedule (Definition 4) and $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ is confluent, then $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{llvm}},\mathcal{C}_{\mathrm{llvm}},\to_{\mathrm{llvm}})$ are equivalent (Definition 2).
Proof Sketch.
Assume $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{llvm}},\mathcal{C}_{\mathrm{llvm}},\to_{\mathrm{llvm}})$ are equivalent on some canonical schedule $\mathsf{Canon}$ and $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ is confluent. Let $c_{\mathrm{df}}\sim_{\mathcal{I}}c_{\mathrm{seq}}$ be any pair of corresponding initial states with $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$, and let $\tau_{\mathrm{seq}}=c_{\mathrm{seq}}\dots$ be the unique complete trace. Let $\tau^{\prime}_{\mathrm{df}}=c_{\mathrm{df}}\dots$ be any complete trace from $c_{\mathrm{df}}$. We need to show that $\tau^{\prime}_{\mathrm{df}}$ and $\tau_{\mathrm{seq}}$ are memory-synchronizing.
By equivalence on $\mathsf{Canon}$, we know that $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ is a complete trace and $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ and $\tau_{\mathrm{seq}}$ are memory-synchronizing. We have two cases:
  • If $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ is finite, then by confluence, $\tau^{\prime}_{\mathrm{df}}$ is finite and has the same final state as $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ (since $\tau^{\prime}_{\mathrm{df}}$ is complete).
  • If $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ is infinite, then by confluence, $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ and $\tau^{\prime}_{\mathrm{df}}$ synchronize infinitely often.
Therefore $\tau^{\prime}_{\mathrm{df}}$ and $\tau_{\mathrm{seq}}$ are memory-synchronizing. ∎
@@
In this section, we present a technique using what we call linear permission tokens to ensure that the dataflow program is confluent. Linear permission tokens are, in essence, a dynamic variant of fractional permissions (Boyland, 2013). In our case, the exact placement of permission tokens is resolved dynamically via symbolic execution.
@@
We associate each value flowing through a dataflow program with a permission token in the form of $\texttt{write}\ l$, $\texttt{read}\ l$, or any formal sum of them, for any memory location $l$. These tokens are used by dataflow operators in a linear fashion: they cannot be replicated or generated, and they are disjoint in the initial state. To support independent parallel reads and enforce that writes must be exclusive, a write token can be split into $k$ many read tokens (for a predetermined parameter $k$), and inversely, $k$ of such read tokens can be merged to get back a write token.
We show in this section that if there is a consistent assignment of these linear permission tokens to values, then we can conclude that the execution of dataflow program will be confluent. This is done in two steps:
@@
  • In Section 5.1, we prove that the technique works in a bounded setting for a concrete trace $\tau$: if $\tau$ has a consistent assignment of permission tokens, $\tau^{\prime}$ is another trace from the same initial configuration, and $\tau^{\prime}$ is bounded by $\tau$ (meaning that for each operator $o$, the number of times $o$ fired in $\tau^{\prime}$ is less than or equal to the number of times $o$ fired in $\tau$), then there is a valid trace from the final configuration of $\tau^{\prime}$ to the final configuration of $\tau$; i.e., $\tau^{\prime}$ will converge to the final configuration of $\tau$.
@@
  • In Section 5.2, we explain how this bounded confluence check can be extended to a full confluence check for the entire dataflow program by using the cut points that we have placed for the dataflow program in Section 4.
Finally in Section 5.3, we show that even though we assume the channels in a dataflow program are unbounded, the cut-simulation and confluence checks allow us to extend the verification results to a bounded-channel model of dataflow programs.
@@ 5.1. Confluence on Bounded Traces
In this section, we show that if we have a finite trace $c_{1}\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}c_{n+1}$ (where $o_{1},\dots,o_{n}$ are the operators fired in each step) with a consistent assignment of permission tokens to each value in each configuration, then for any other trace $c_{1}\to^{o^{\prime}_{1}}_{\mathrm{df}}\dots\to^{o^{\prime}_{m}}_{\mathrm{df}}c_{m+1}$ (not necessarily with a consistent assignment of permission tokens) such that the multiset $\{o^{\prime}_{1},\dots,o^{\prime}_{m}\}\subseteq\{o_{1},\dots,o_{n}\}$, we have $c_{m+1}\to^{*}_{\mathrm{df}}c_{n+1}$.
@@
  • $P(L,k)$ is the set of permissions such that $(P(L,k),0,+)$ forms a partially commutative monoid. It consists of formal sums $\texttt{read}\ l_{1}+\dots+\texttt{read}\ l_{n}$, with $l_{1},\dots,l_{n}\in L$, modulo the commutativity and associativity of $+$. We require that for a location $l$, at most $k$ copies of $\texttt{read}\ l$ are present in the formal sum. We denote $N(l,p)$ to be the number of occurrences of $\texttt{read}\ l$ in $p$. We call $p+q$ the (disjoint) sum of $p,q\in P(L,k)$. In particular, for any $l\in L$, $p+q$ is defined iff for all $l$, $N(l,p)+N(l,q)\leq k$.
    We also denote
@@
Now we define when two or more permissions are considered disjoint, which is crucial for enforcing linearity.
Definition 0.
[…]
For example, in $P(\{\texttt{A},\texttt{B}\},2)$, we have that
@@
Intuitively, the reason to allow a write to split into $k$ copies of read is to allow the following two valid scenarios:
  • A store operator can write to a memory location if it has exclusive write permission to the location (i.e., no other operator has a write or read permission to that memory location);
  • A load operator can read from a memory location if it has a read permission and all other currently existing permissions in the system are also read.
Furthermore, once (at most $k$) parallel reads are performed, a store operator can perform a memory write by reclaiming all existing read permissions and merging them into a write; hence we define write as the sum of $k$ copies of read.
Now let us define the assignment of permission tokens to values in a configuration. Let $E$ denote the set of channel names. For a configuration $c\in\mathcal{C}_{\mathrm{df}}$ and a channel $e\in E$, let $\mathrm{Channel}(c,e)$ denote the state of the channel $e$ in $c$, which is a string over $V$ representing the values in the channel $e$.
@@
Let $c\in\mathcal{C}_{\mathrm{df}}$ be a dataflow configuration. A permission augmentation of $c$ is a partial map $t:E\times\mathbb{N}\rightharpoonup P(L,k)$ such that for any $e\in E$, $t(e,n)$ is defined iff $n\in\{1,\dots,|\mathrm{Channel}(c,e)|\}$. We call $(c,t)$ a permission-augmented configuration.
Intuitively, $t(e,n)$ is the permission we attach to the $n$-th value in the channel $e$. Permission augmentations defined above can assign arbitrary permissions to values. However, to ensure race-freedom and confluence, we need to restrict them so that, for example, two operators cannot share a write token. In the following definitions, we define when a configuration, a transition, and a trace are consistently augmented with permissions, meaning that permissions are used linearly.
@@
Let $c\to^{o}_{\mathrm{df}}c^{\prime}$ be a transition. Let $t,t^{\prime}$ be two consistent permission augmentations for $c,c^{\prime}$, respectively. Let $p_{1},\dots,p_{n}$ be the permissions in $t$ attached to input values of $o$ and $q_{1},\dots,q_{m}$ be the permissions in $t^{\prime}$ attached to output values of $o$. We say $(c,t)\to^{o}_{\mathrm{df}}(c^{\prime},t^{\prime})$ is a consistent (permission-augmented) transition iff:
  • $(c,t)$ and $(c^{\prime},t^{\prime})$ are consistent.
  • $t(e,n)=t^{\prime}(e,n)$ for all channel $e$ and position $n$ except for those changed by $o$.
@@
Let $c_{1}\to_{\mathrm{df}}\dots\to_{\mathrm{df}}c_{n}$ be a trace. Let $t_{1},\dots,t_{n}$ be permission augmentations to $c_{1},\dots,c_{n}$, respectively. We say that $(c_{1},t_{1})\to_{\mathrm{df}}\dots\to_{\mathrm{df}}(c_{n},t_{n})$ is a consistent (permission-augmented) trace iff for all $i\in\{1,\dots,n\}$, $(c_{i},t_{i})\to_{\mathrm{df}}(c_{i+1},t_{i+1})$ is consistent.
We now show some facts about a consistently permission-augmented transition or trace, leading to the final Theorem 11. First, we show that two consistent steps commute.
Lemma 9.
@@
Proof Sketch.
If one of $o_{1}$ and $o_{2}$ is neither load nor store, or if both of them are loads, they can trivially commute with the same result. Otherwise, one of $o_{1}$ and $o_{2}$ is a store operator (without loss of generality, assume $o_{1}$ is a store). For $i\in\{1,2\}$, let $l_{i}$ be the memory location accessed by $o_{i}$ and $p_{i}$ be the sum of $o_{i}$’s input permissions. Since they can both fire at $c_{1}$, they should have suitable permissions: $\texttt{write}\ l_{1}\leq p_{1}$ and $\texttt{read}\ l_{2}\leq p_{2}$. Since $p_{1}$ and $p_{2}$ are disjoint, we have $l_{1}\neq l_{2}$ Therefore, they are accessing different memory locations, and should give the same result no matter in which order we execute them. ∎
@@
Although we have omitted much of the technical details of the proof, we have formalized a machine-checkable proof of Theorem 11 in Verus (Lattuada et al., 2023), a verification language, and the proof can be found in our GitHub repository (Lin et al., 2024).
Practically, with this bounded confluence theorem, we can do bounded model checking for the confluence property: we perform symbolic execution to obtain a trace of symbolic configurations $c_{1}(\mathbf{x})\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}c_{n+1}(\mathbf{x})$, and then we can solve for a consistent augmentation $(t_{1},\dots,t_{n+1})$ for this trace using an SMT solver. If there is a solution, then it follows from Theorem 11 that any other trace (firing operators only in the multiset $\{o_{1},\dots,o_{n}\}$, counting multiplicity) would still converge to $c_{n+1}(\mathbf{x})$ in the end.
Algorithm 1 Full confluence checking algorithm (Section 5.2).
1: Inputs:
  • $c^{1}_{\mathrm{df}}(\mathbf{x}^{1}),\dots,c^{n}_{\mathrm{df}}(\mathbf{x}^{n})$: a list of dataflow cut points;
  • $P(L,k)$: a permission algebra.
2:If ConfluenceCheck succeeds, then the dataflow program is confluent.
3:procedure ConfluenceCheck
4:     $\psi\leftarrow\top$triangleright Permission constraints to be accumulated.
5:     for $i\in\{1,\dots,n\}$ do $\triangleright$ Initialize permission variables at cut points
[…]
7:         for each channel $e$ and $m\in\{1,\dots,|\mathrm{Channel}(c^{i}_{\mathrm{df}}(\mathbf{x}^{i}),e)|\}$ do
[…]
10:     for $i\in\{1,\dots,n\}$ do
[…]
12:         $\textbf{enqueue}(Q,(c^{i}_{\mathrm{df}}(\mathbf{x}^{i}),t_{i}))$
13:         while $Q$ is non-empty do
14:              $(c_{\mathrm{df}}(\mathbf{x}^{i}),t)\leftarrow\textbf{dequeue}(Q)$
15:              for $(c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i}),t^{\prime})\in\textbf{step\textunderscore canonical}(c_{\mathrm{df}}(\mathbf{x}^{i}),t)$ do $\triangleright$ For each symbolic branch
16:
17:                  $\triangleright$ Constraints in Definition 7, where $o$ is the fired operator.
18:                  $p_{\mathrm{in}}\leftarrow$ sum of $o$’s input permission variables in $t$
19:                  $p_{\mathrm{out}}\leftarrow$ sum of $o$’s output permission variables in $t^{\prime}$
20:                  $\psi\leftarrow\psi\wedge(p_{\mathrm{out}}\leq p_{\mathrm{in}})$
21:                  if $o$ reads memory location $l\in L$ then
22:                       $\psi\leftarrow\psi\wedge(\texttt{read}\ l\leq p_{\mathrm{in}})$
23:                  else if $o$ writes to memory location $l\in L$ then
24:                       $\psi\leftarrow\psi\wedge(\texttt{write}\ l\leq p_{\mathrm{in}})$
25:
26:                  for $j\in\{1,\dots,n\}$ do
27:                       if $\exists\sigma:\mathbf{x}^{j}\to V(\mathbf{x}^{i}).\ c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i})=c^{j}_{\mathrm{df}}(\sigma)$ then $\triangleright$ Matched to a cut point
28:                           for each channel $e$ and $m\in\{1,\dots,|\mathrm{Channel}(c^{j}_{\mathrm{df}}(\mathbf{x}^{j}),e)|\}$ do
29:                                $\psi\leftarrow\psi\wedge(t^{\prime}(e,m)=t_{j}(e,m))$
30:                           continue Line 15
31:                  $\textbf{enqueue}(Q,(c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i}),t^{\prime}))$
32:     succeed iff $\psi$ is satisfiable in $P(L,k)$
@@ 5.2. Full Confluence Checking using Cut Points
To extend the bounded confluence results in Section 5.1 to checking confluence for the entire dataflow program with unbounded traces, we need to finitely describe and check a consistent pattern of permission augmentations for all traces. For this purpose, we use the cut points selected in Section 4 and perform symbolic execution from them not only for symbolic values, but also for permissions.
@@
Let $P^{1},\dots,P^{n}$ be a list of cut points from the cut-simulation check (Section 4.3), where each $P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x}^{i}),c^{i}_{\mathrm{seq}}(\mathbf{y}^{i}))\mid\varphi^{i}(\mathbf{x}^{i},\mathbf{y}^{i})$ is a pair of symbolic dataflow and Seq configurations with an additional correspondence constraint. During the cut-simulation check in Section 4.3, we verify that starting from any dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x}^{i})$, any symbolic execution branch reaches an instance of another cut point $c^{j}_{\mathrm{df}}(\mathbf{x}^{j})$. A consequence is that the symbolic branches from all dataflow cut points “cover” any complete trace in the canonical schedule (Definition 3). In other words, any complete concrete trace in the canonical schedule is the concatenation of instances of symbolic branches.
@@
Thus, if we are able to find a consistent permission augmentation for the trace of each dataflow symbolic branch during the cut-simulation check, then any concrete trace in the canonical schedule would have a consistent permission augmentation, which implies that any other schedule will converge to the canonical schedule, and the confluence of the entire dataflow program follows.
This full confluence check algorithm is procedurally described in Algorithm 1. The algorithm is parameterized with a list of dataflow cut points $c^{1}_{\mathrm{df}}(\mathbf{x}^{1}),\dots,c^{n}_{\mathrm{df}}(\mathbf{x}^{n})$ and a predetermined permission algebra $P(L,k)$. The algorithm first generates a fresh permission variable for each value present in the channels of the dataflow cut points. Then it performs symbolic execution from each cut point, carrying constraints for both symbolic values and their corresponding permission variables, until the symbolic branch is subsumed by another cut point. In the end, it checks for the satisfiability of the permission constraints. If they are satisfiable, the confluence of the full program follows.
In more detail, the first loop at Line 5 assigns fresh permission variables to each value present in each cut point, representing an unknown concrete permission token to be solved for. We also add disjointness constraints at Line 9 to enforce that these permission augmentations are consistent (Definition 6) at the corresponding cut point configurations.
Then in the main loop at Line 10, we symbolically execute from each dataflow cut point, trying to match each symbolic branch against other cut points. The step function step_canonical uses a deterministic canonical schedule (Definition 3) we constructed from Seq execution traces. For example, in the case of LLVM and RipTide, we use the canonical schedule described in Section 4. step_canonical is also instrumented to take a permission augmentation $t$ for the current configuration $c_{\mathrm{df}}(\mathbf{x}^{i})$, and output a modified permission augmentation $t^{\prime}$ for the stepped configuration $c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i})$, in which fresh variables are generated for all new output values in $c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i})$, and permissions corresponding to input values are removed. In each step, for each symbolic branch, we also accumulate permission constraints (Lines 17-24) to enforce that the augmented transition $(c_{\mathrm{df}}(\mathbf{x}^{i}),t)$ to $(c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i}),t^{\prime})$ is a consistent transition (Definition 7).
At each symbolic step, we check if a symbolic branch is matched to another cut point $c^{j}_{\mathrm{df}}(\mathbf{x}^{j})$ (Line 27). This check is performed by finding a substitution $\sigma:\mathbf{y}^{j}\to\mathsf{Terms}(\mathbf{x}^{i})$ such that the current symbolic branch $c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i})=c^{j}_{\mathrm{df}}(\sigma)$. If the symbolic branch matches the cut point, we add equality constraints (Line 29) to enforce that the permissions in $t^{\prime}$ should converge back to the original permissions $t_{j}$ at the matched $j$-th cut point. Note that the substitution $\sigma$ is not later used because we do not need to check correspondence constraints (Section 4.3).
The main loop at Line 10 should terminate, because if the cut-simulation check (Section 4) passes, any branch from any dataflow cut point should eventually match another cut point.
Finally, we check if the accumulated permission constraint $\psi$ is satisfiable in the permission algebra $P(L,k)$ via an SMT query. If $\psi$ is satisfiable, then any concrete trace (of unbounded length) in the canonical schedule has a consistent augmentation. Thus, by Theorem 11, any other possible schedule would eventually converge to the canonical schedule, and the dataflow program is confluent. If $\psi$ is not satisfiable, then there are different possibilities: 1) the dataflow program is not confluent and may have data races (which require manual inspection to confirm); 2) the choice of $k$ in the permission algebra $P(L,k)$ is not large enough; or 3) the confluence property is not provable using this method.
5.3. Liveness in the Bounded-Channel Model
In our current semantics for dataflow programs, channels between operators are modeled as unbounded queues. In real CGRA architectures, however, the channels have a fixed buffer size, and an operator may block if one of the output channels is full. As a result, a dataflow program in the bounded-channel model may have a deadlock that is not possible in the unbounded-channel model.
For example, consider Figure 7, where operator A has two dataflow paths to B: one path is full, while the other is empty due to a Steer operator T discarding its inputs. This situation puts us in deadlock. Operator B cannot fire without tokens along both paths, while operator A cannot fire before more space is made along the full path.
To verify that this scenario does not occur, we lift our cut-simulation and confluence results in the unbounded-channel model to the bounded-channel model via two observations:
  • The symbolic branches in the dataflow cut point execution (Section 4.3) have finitely many intermediate symbolic configurations, so the maximum channel size is bounded by some integer $K$ in the canonical schedule.
  • The confluence results in Section 5.1 and Section 5.2 are still valid in the bounded-channel model, which we have formalized and verified in Verus as well.
Therefore, if the cut-simulation and the confluence checks pass in the unbounded-channel model, the same results hold in the bounded-channel model if channels have a buffer size of at least $K$. As a result, in the bounded-channel model, the dataflow program is still equivalent to the input imperative program, which guarantees the liveness property of the dataflow program that it will always make progress if the sequential imperative program is able to make progress. Furthermore, since the sequential program is deadlock-free, the dataflow program is also deadlock-free.
Name LOC #OP #PC Sim. Conf. Description
nn_vadd 5 15 244 0.02 0.17 Vector addition (RipTide)
nn_norm 10 19 316 0.02 0.19 Neural network normalization (RipTide)
nn_relu 11 19 844 0.03 0.45 Neural network ReLU layer (RipTide)
smv 12 25 1103 0.05 2.27 Sparse-dense matrix-vector mult. (Pipestitch)
dmv 10 31 1234 0.06 1.64 Dense-dense matrix-vector mult. (RipTide)
Dither 19 31 2553 0.13 2.29 Dithering (Pipestitch)
SpSlice 30 31 2616 $\times$ 0.13 7.94 Sparse matrix slicing (Pipestitch)
nn_fc 23 36 2632 0.09 3.35 Neural network fully-connected layer (RipTide)
SpMSpVd 45 40 3182 0.15 8.38 Sparse-sparse matrix-vector mult. (Pipestitch)
bfs 48 48 3634 0.15 7.68 Breadth-first search (RipTide)
dfs 45 49 3762 0.17 7.76 Depth-first search (RipTide)
smm 22 49 3967 0.16 11.2 Sparse-dense matrix mult. (RipTide)
nn_pool 32 52 8165 0.23 7.14 Neural network pooling layer (RipTide)
dmm 16 56 4220 0.19 6.16 Dense-dense matrix mult. (RipTide)
sconv 22 56 4272 0.2 10.37 Sparse convolution (RipTide)
sort 23 57 6554 $\times$ 0.14 3.31 Radix sort (RipTide)
SpMSpMd 39 60 7693 0.33 26.68 Sparse-sparse matrix mult. (Pipestitch)
nn_conv 42 61 12154 0.2 15.32 Neural network convolution layer (RipTide)
dconv 24 65 8134 0.25 11.51 Dense convolution (RipTide)
fft 29 70 2690 0.12 5.77 Fast Fourier transform (RipTide)
sha256 81 181 3548 $\times$ 0.28 12.6 SHA-256 hash